Problem 10-9 (Algo) Comprehensive Variance Analysis [LO10-1, LO10-2, LO10-3] Marvel Parts, Inc., manufactures auto accessories. One of the company’s products is a set of seat covers that can be adjusted to fit nearly any small car. The company has a standard cost system in use for all of its products. According to the standards that have been set for the seat covers, the factory should work 990 hours each month to produce 1,980 sets of covers. The standard costs associated with this level of production are: Total Per Set of Covers Direct materials $ 45,738 $ 23.10 Direct labor $ 6,930 3.50 Variable manufacturing overhead (based on direct labor-hours) $ 3,168 1.60 $ 28.20 During August, the factory worked only 1,000 direct labor-hours and produced 2,500 sets of covers. The following actual costs were recorded during the month: Total Per Set of Covers Direct materials (10,000 yards) $ 56,000 $ 22.40 Direct labor $ 9,250 3.70 Variable manufacturing overhead $ 4,500 1.80 $ 27.90 At standard, each set of covers should require 3.3 yards of material. All of the materials purchased during the month were used in production. Required: 1. Compute the materials price and quantity variances for August. 2. Compute the labor rate and efficiency variances for August. 3. Compute the variable overhead rate and efficiency variances for August.
Answer #1
Standard materials price per yard 7.00 =23.1/3.3
Standard labor hour per unit 0.5 =990/1980
Standard labor rate per labor hour 7 =3.5/0.5
Standard variable rate per labor hour 3.2 =1.6/0.5
1
Materials price variance 14000 F =56000-(10000*7)
Materials quantity variance 12250 U =7*(10000-2500*3.3)
2
Labor rate variance 2250 U =9250-(1000*7)
Labor efficiency variance 1750 F =7*(1000-2500*0.5)
3
Variable overhead rate variance 1300 U =4500-(1000*3.2)
Variable overhead efficiency variance 800 F =3.2*(1000-2500*0.5)